# User Simulation for Evaluating Information Access Systems

### v1

@@ Abstract
This book focuses on providing a thorough understanding of user simulation techniques designed specifically for evaluation purposes. We begin with a background of information access system evaluation and explore the diverse applications of user simulation. Subsequently, we systematically review the major research progress in user simulation, covering both general frameworks for designing user simulators, utilizing user simulation for evaluation, and specific models and algorithms for simulating user interactions with search engines, recommender systems, and conversational assistants. Realizing that user simulation is an interdisciplinary research topic, whenever possible, we attempt to establish connections with related fields, including machine learning, dialogue systems, user modeling, and economics. We end the book with a detailed discussion of important future research directions, many of which extend beyond the evaluation of information access systems and are expected to have broader impact on how to evaluate interactive intelligent systems in general.
@@
Information access systems, such as search engines, recommender systems, and conversational assistants, have become increasingly intelligent, thanks to recent progress in research in artificial intelligence (AI), especially in machine learning and natural language processing. These information access systems are used by millions on a daily basis and focus on tasks where humans really need help: finding information. In general, the interactions with these systems involve entering information needs or preferences (by typing queries, rating items, or asking natural language questions) and interacting with information objects (by clicking, typing, or speaking) that are presented by the system on some device (e.g., desktop, laptop, tablet, smart phone, or smart speaker) in some modality or combination of modalities (e.g,. text, rich snippets, voice). With intelligent home devices becoming available, information access systems may also be used to power a wide range of intelligent agent systems, such as Google Home or Alexa, which can go beyond supporting information access to also support other user tasks (e.g., controlling home appliances, making appointments, or placing orders).
@@ 1.1 Information Access Tasks
Information access refers to the ability to identify, retrieve, and use information effectively. [^1] https://www.encyclopedia.com/computing/news-wires-white-papers-and-books/information-access Access to the right information at the right time plays an important role in everyone’s life and is vital to business operations. At a high-level, information access can happen in two modes [Zhai:2016:MC]: (1) pull mode, where the user takes the initiative and uses a search engine to find needed information (“pull" relevant information to the user), and (2) push mode, where the system takes the initiative and recommends relevant information to a user (“push" relevant information to the user). Search engines and recommender systems are the two most common widely used applications for information access. The two modes of information access are complementary and often supported simultaneously using a single system. For example, a search engine can not only support querying (pull mode) but also recommend related information to the user (push mode). Similarly, a recommender system may also recommend information in the form of a ranked list to enable a user to further interact with the recommended information and potentially enter a query to further explore the information space. Indeed, search and recommendation have been suggested as “two sides of the same coin” in [Belkin:1992:CACM]. As such, it is not surprising that search engines and recommender systems share many common technical challenges (e.g., modeling a user’s information need and preferences, matching an information item with a user’s interest, ranking items accurately, learning from a user’s feedback information, evaluating a ranked list to assess its utility to a user), and tend to benefit from using similar techniques, including user simulation techniques. For this reason, we intend to cover the topic of user simulation in the broad context of information access with the understanding that most discussions are relevant to both search engines and recommender systems, even though the actual research work that we discuss may have been done for either just search engines or recommender systems.
@@
Recently, conversational assistants have attracted much attention. They generally support mixed-initiative interaction via natural language to facilitate both search and recommendation in the same information access session [Zamani:2023:arXiv]. Compared with traditional search engines and recommender systems, where the actions a user could potentially take are well specified by the user interface of the system, conversational assistants have more open-ended functions in the sense that a user can potentially ask questions, provide clarifications, and explore related topics using unrestricted natural language, thus adding complexity to user simulation. We note that conversational assistants, casually referred to as “conversational AI,” can cater to diverse user goals, including, e.g., social chatting. However, our primary focus in this book is on systems that are designed to support information access, i.e., tasks where there is an underlying information need and the system returns information objects (which may be documents, entities, answers, utterances, etc.). This focus increases the commonality between conversational assistants, search engines, and recommender systems in terms of user simulation.
1.2 Evaluation Goals and Methodologies
1.2.1 Evaluation Goals
We may distinguish between two main evaluation goals, which have different requirements:
  • To know the actual utility of technology, that is, how useful is a given system to a group of users? This question is important for assessing the potential of building a product based on the technology. One needs to answer such a question with an interpretable absolute value to quantify the performance and utility of the system, which can then be used to inform decisions like: “Is it worth the investment?” or “Is it worth deploying?”
  • To establish a relative comparison between two systems, which is required in order to make progress in developing better systems and advance research. This is a weaker requirement that only needs a relative measure that is correlated with the absolute difference. At the same time, reproducibility is a necessary condition.
1.2.2 Evaluation Methodologies
@@
Reusable test collections facilitate large-scale automatic evaluation and have been invaluable for comparing different algorithms and improving the state of the art. They ensure repeatability and enable comparison between different approaches and study of effectiveness of individual components in a complex method. However, they are static and are severely limited in their ability to capture many aspects of users and interactions adequately. They measure system performance on an abstraction of a given process (e.g., search or recommendation), where the user is also abstracted away. This has so far been the standard evaluation methodology for making relative comparisons between two systems in a repeatable and reproducible manner. However, it is generally not so useful for the purpose of evaluating the actual utility of a system due to the significant deviation of the evaluation environment from the real application environment, especially due to the very limited inclusion of users.
@@
User studies provide the highest fidelity in terms of capturing real users’ interactions with an actual system in a controlled setting. However, experiments that involve real users are costly to run. Further, if multiple rounds of experimentation are performed, new users need to be involved in order to avoid misinterpretation of the findings due to fatigue and learning effects. Also, experiments that involve an actual service have a bandwidth limit, which is set by the amount of users and their activity using that service. User studies are useful for assessing the actual utility of a system, but they suffer from the limitation of being not reproducible (even the same user would not behave in the same way when repeating an experiment due to learning effects). Thus, they have limited value for making relative comparisons between systems, especially when new systems—to be developed in the future—need to be included in the comparison.
@@
Online evaluation (a.k.a. log-based studies) is based on the idea of observing real users of a fully operational system and assessing the system’s performance by analyzing the recorded user behaviour. It is widely used by companies that deploy real-world applications or services, and is regarded as the most direct and reliable measurement of quality and user experience. Like user studies, online evaluation enables measuring the actual utility of a system and comparing systems with real users, but at a much larger scale in terms of the number of users used for evaluation. Unlike user studies, however, it generally does not provide control over the users; also, users’ interaction are limited to natural interactions with the system, making it harder to interpret the results. As in the case of user studies, online evaluation suffers from being not reproducible and thus cannot be “reused” to compare different systems or analyze the effectiveness of various components.
@@ 1.3 Challenges in Evaluating Information Access Systems and Simulation-based Evaluation
In general, all the three methodologies discussed earlier can be applied to, and indeed have been regularly used for, evaluating information access systems. However, none of those methodologies can be used to compare multiple interactive information access systems (in terms of their overall effectiveness in supporting users) using reproducible experiments due to the static nature of the test collection-based approach and the lack of reproducibility when real users are involved.
It is important to evaluate the overall effectiveness of a system, especially in the case of information access systems. This is because users are likely to benefit the most from intelligent assistance in case of complex tasks. Commonly, complex tasks are decomposed into a series of smaller and simpler components. The decomposition process generally involves close collaboration between a user and a system in an interactive way, in which a user would iteratively direct the system to perform specific component functions and the results from multiple steps can be synthesized to generate solutions to a complex problem. Most component-level tasks can be abstracted, studied and addressed in isolation, and evaluated using the reusable test collection methodology with reproducible experiments. While this is clearly necessary to allow for systematic progress to be made, evaluation of individual components alone is insufficient. It is arguably more important to view them as components of a larger information access system and study how to evaluate the whole system from a user’s perspective. Indeed, the ultimate goal of evaluation is to measure how well the user is aided in achieving their end goal. It is vitally important to do this evaluation correctly; if not done appropriately, it would mislead us to draw wrong conclusions or deploy an inferior application system, which hurts users.
Logically, the evaluation of an interactive system’s overall effectiveness must involve a user in some way. The fact that the involvement of real users inherently leads to non-reproducible experiments suggests the need for simulating real users. Simulated users can be controlled and thus enable reproducible experiments. If a user simulator can simulate a real user well, such a user simulation-based evaluation method would solve the challenge of evaluating interactive information access systems with reproducible experiments.
@@ 1.4 User Simulation
What do we exactly mean by user simulation? Informally, having an intelligent agent to simulate how a user interacts with a system. The agent can be built based on models/algorithms/rules and any knowledge we have about the user (their behaviour, knowledge, etc.). The agent can also have parameters that can be varied in a meaningful way to simulate variations of users. Once a user simulator is constructed, it can then be used to interact with any system that needs to be evaluated. In turn, we can measure the system’s utility based on the observed behaviour and interaction data of the agent while interacting with the system. For relative comparisons of systems, simulation does not need to be perfect; it is enough to identify relative system differences.
Simulation has the potential to enable repeatable and reproducible evaluations at a low cost, without using invaluable user time (human assessor time or online experimentation bandwidth). Further, simulation can augment traditional evaluation methodologies by offering possibilities to gain insights into how system performance changes under different conditions and user behaviour. In general, user simulation has many uses, including at least the following:
@@
  • Generating synthetic data with the purpose of training machine learning models, especially reinforcement learning.
We note that all these uses require similar techniques, but our focus is in evaluation. One important difference is that for that purpose, simulation does not need to be perfect; it is enough to identify relative system differences. For example, although the relevance judgments in almost all the test collections for information retrieval evaluation are incomplete, the conclusions about relative performance of different retrieval systems tend not be affected much by the approximation made in a test collection [Voorhees:2000:IPM].
1.4.1 Simulation Approaches
In general, all approaches to simulation of users assume that there is initially some information available, including (1) information about the system that user interacts with (e.g., a search engine) and its user interface (e.g., a query box and navigable search result lists), (2) information about the user’s task (e.g., collecting as many relevant information items as possible or finding a suitable product to purchase), and (3) any other potentially useful information about the user (e.g., background knowledge, context). Those approaches would then attempt to simulate all the actions a user can potentially take when interacting with the system to perform the task, given any particular interaction context. Thus, approaches can be classified based on the specific actions that they attempt to simulate. For example, some approaches may simulate how a user generates a query while others might simulate how a user responds to a search result list (e.g., simulating when a user might click on a result and when the user might skip one). Approaches simulating different types of actions can also be combined to simulate a whole session of actions of a user. As will be elaborated later in the book, most existing work tends to focus on simulating each type of actions separately with significantly less work on simulating a whole user session.
The problem of simulating a user’s action can often be framed as a classification problem when there are a finite number of actions to choose from; for example, simulation of a user’s clicking action may be framed as a binary classification problem, where the algorithm would predict whether the simulated user would click on a result or not after examining a snippet. When there are potentially infinitely many actions to choose from (e.g., when formulating a query, any valid query would be potentially an option), in practice, we often make assumptions to restrict the number of actions to be considered when simulating those actions (e.g., limit the length of a query to be considered).
With the problem framed as a classification problem, different approaches generally vary in how they perform the classification (equivalently prediction) task. At a high level, we can distinguish two broad approaches: model-based and data-driven. Model-based approaches may be based on rules designed based on knowledge about how users behave or on interpretable probablistic models that can more flexibly capture uncertainties. The parameters of such models may be heuristically set or empirically estimated based on observed user data. Varying those parameters often also leads to simulation of different kind of users. Data-driven approaches emphasize maximizing accuracy of fitting any observed real user data, without necessarily imposing interpretability. Almost all such approaches are based on supervised machine learning, notably using deep neural networks which can learn effective, but non-interpretable representations from the data for predictive modeling. The two families of approaches may also be combined, e.g., by utilizing model-based techniques to compute effective features for data-driven approaches. However, interpretability is desirable when building user simulators for evaluation to ensure that evaluation results are meaningful. Hence, this book primarily focuses on interpretable model-based approaches, enabling the testing of verifiable hypotheses about users and facilitating interpretable evaluations through meaningful variations in user simulation.

### v2

@@ Abstract
This book focuses on providing a thorough understanding of user simulation techniques designed specifically for evaluation purposes. We begin with a background of information access system evaluation and explore the diverse applications of user simulation. Subsequently, we systematically review the major research progress in user simulation, covering both general frameworks for designing user simulators, utilizing user simulation for evaluation, and specific models and algorithms for simulating user interactions with search engines, recommender systems, and conversational assistants. Realizing that user simulation is an interdisciplinary research topic, whenever possible, we attempt to establish connections with related fields, including machine learning, dialogue systems, user modeling, and economics. We end the book with a broad discussion of important future research directions, many of which extend beyond the evaluation of information access systems and are expected to have broader impact on how to evaluate interactive intelligent systems in general.
@@
Information access systems, such as search engines, recommender systems, and conversational assistants, have become increasingly intelligent in understanding users’ intents, supporting their tasks, and interacting with them using natural language dialogue, thanks to recent progress in research in artificial intelligence (AI), especially in machine learning and natural language processing. These information access systems are used by millions on a daily basis to perform a wide range of tasks where humans need help to find information relevant to a task. In general, the interactions with these systems involve a user entering information needs or preferences (by typing queries, rating items, or asking natural language questions) and interacting with information objects (by clicking, typing, or speaking) that are presented by the system on some device (e.g., desktop, laptop, tablet, smart phone, or smart speaker) in some modality or combination of modalities (e.g,. text, rich snippets, voice). With intelligent home devices becoming available, information access systems may also be used to power a wide range of intelligent agent systems, such as Google Home or Alexa, which can go beyond supporting information access to also support other user tasks (e.g., controlling home appliances, making appointments, or placing orders).
@@ 1.1 Information Access Tasks
Information access refers to the ability to identify, retrieve, and use information effectively. [^1] https://www.encyclopedia.com/computing/news-wires-white-papers-and-books/information-access Access to the right information at the right time plays an important role in everyone’s life and is vital to business operations. At a high-level, information access can happen in two modes [Zhai:2016:MC]: (1) pull mode, where the user takes the initiative and uses a search engine to find needed information (“pull" relevant information to the user), and (2) push mode, where the system takes the initiative and recommends relevant information to a user (“push" relevant information to the user). The two modes can be naturally mixed in conversational AI systems, which are increasingly common due to the emergence of large language models (LLMs) [McTear:2024:book].
@@
Search engines and recommender systems are the two most common widely used applications for information access. The two modes of information access are complementary and often supported simultaneously using a single system. For example, a search engine can not only support querying (pull mode) but also recommend related information to the user (push mode). Similarly, a recommender system may also recommend information in the form of a ranked list to enable a user to further interact with the recommended information and potentially enter a query to further explore the information space. Indeed, search and recommendation have been suggested as “two sides of the same coin” in [Belkin:1992:CACM]. As such, it is not surprising that search engines and recommender systems share many common technical challenges (e.g., modeling a user’s information need and preferences, matching an information item with a user’s interest, ranking items accurately, learning from a user’s feedback information, evaluating a ranked list to assess its utility to a user), and tend to benefit from using similar techniques, including user simulation techniques. For this reason, we intend to cover the topic of user simulation in the broad context of information access with the understanding that most discussions are relevant to both search engines and recommender systems, even though the actual research work that we discuss may have been done for either just search engines or recommender systems.
@@
Recently, conversational assistants have attracted much attention [McTear:2021:book]. They generally support mixed-initiative interaction via natural language to facilitate both search and recommendation in the same information access session [Zamani:2023:FnTIR]. Compared with traditional search engines and recommender systems, where the actions a user could potentially take are well specified by the user interface of the system, conversational assistants have more open-ended functions in the sense that a user can potentially ask questions, provide clarifications, and explore related topics using unrestricted natural language, thus adding complexity to user simulation. We note that conversational assistants, casually referred to as “conversational AI,” can cater to diverse user goals, including, e.g., social chatting. However, our primary focus in this book is on systems that are designed to support information access, i.e., tasks where there is an underlying information need and the system returns information objects (which may be documents, entities, answers, utterances, etc.). This focus increases the commonality between conversational assistants, search engines, and recommender systems in terms of user simulation.
1.2 Evaluation Methodologies
@@
Reusable test collections (a.k.a. offline evaluation) facilitate large-scale automatic evaluation and have been invaluable for comparing different algorithms and improving the state of the art. They ensure repeatability and enable comparison between different approaches and study of effectiveness of individual components within complex methods. However, they are static and are severely limited in their ability to capture many aspects of users and interactions adequately. They measure system performance on an abstraction of a given process (e.g., search or recommendation), where the user is also abstracted away. Specifically, they are based on simplified models of the information access process and user behaviour. Examples include the assumption that a system always presents a ranked list of results to a user or that a user can always recognize whether a document is relevant in a search result list. This has so far been the standard evaluation methodology for making relative comparisons between two systems in a repeatable and reproducible manner. However, it is generally not so useful for the purpose of evaluating the actual utility of a system due to the significant deviation between the evaluation environment and the real-world application and the very limited inclusion of users. It is also generally hard or impossible to develop test collections for evaluating an interactive system, a limitation that can be addressed by user simulation.
@@
User studies provide the highest fidelity in terms of capturing real users’ interactions with an actual system in a controlled setting. However, experiments that involve real users are costly to run. Further, if multiple rounds of experimentation are performed, new users need to be involved in order to avoid misinterpretation of the findings due to fatigue and learning effects. Also, experiments that involve an actual service have a bandwidth limit, which is set by the amount of users and their activity using that service. User studies are useful for assessing the actual utility of a system, but they suffer from several limitations. First, the result is generally not reproducible (even the same user would not behave in the same way when repeating an experiment due to learning effects). Thus, they have limited value for making relative comparisons between systems, especially when new systems—to be developed in the future—need to be included in the comparison. Second, the cost of recruitment is often high, and it is a challenge to recruit enough people from the right population. Major industry labs often have to invest significantly in recruiting users and conducting user studies, which smaller companies typically cannot afford. In academia, only a few examples of user studies of reasonable scale involve participants beyond university students [Brennan:2014:IIIX].
@@
Online evaluation (a.k.a. log-based studies) is based on the idea of observing real users of a fully operational system and assessing the system’s performance by analyzing the recorded user behaviour. For instance, A/B tests can be used to evaluate different versions of a system. Online evaluation is widely used by companies that deploy real-world applications or services, and is regarded as the most direct and reliable measurement of quality and user experience. Like user studies, online evaluation enables measuring the actual utility of a system and comparing systems with real users, but at a much larger scale in terms of the number of users used for evaluation. Unlike user studies, however, it generally does not provide control over the users, making it harder to interpret the results. As in the case of user studies, online evaluation suffers from being not reproducible and thus cannot be “reused” to compare different systems or analyze the effectiveness of various components. Another limitation of online evaluation is that the user interactions to be evaluated are limited to natural interactions with the system, thus it cannot accommodate counterfactual evaluation, i.e., where the potential outcomes of a hypothetical system (e.g., one with a new algorithm) and being compared to those of an existing system. This hypothetical system cannot be used by real users, making online evaluation infeasible. Additionally, there is a risk of leaving a negative impression on users about a production system’s performance if a system to be evaluated online turns out to perform poorly.
@@ 1.3 Challenges in Evaluating Information Access Systems and Simulation-Based Evaluation
In general, all the three methodologies discussed earlier can be applied to, and indeed have been regularly used for, evaluating information access systems. However, none of those methodologies can be used to compare multiple interactive information access systems (in terms of their overall effectiveness in supporting users) using reproducible experiments due to the static nature of the test collection-based approach and the lack of reproducibility when real users are involved. These limitations can be addressed by using user simulation: simulated users can be controlled and thus enable reproducible experiments.
Note that the existing test collection-based evaluation methodology [Sanderson:2010:FnTIR] can be viewed as a simple form of user simulation. This means we are already utilizing user simulation, albeit implicitly, without explicitly articulating what kind of users are being simulated. A key advantage of evaluation based on user simulation is to make assumptions about simulated users and their behaviour more explicit, while modeling a broader range of user actions than current measures consider (see Section 4.3 for more discussion on this).
We want to highlight the importance of evaluating the overall effectiveness of a system (not just various components of a system), especially in the case of information access systems. This is because users are likely to benefit the most from intelligent assistance in case of complex tasks. Commonly, complex tasks are decomposed into a series of smaller and simpler components. The decomposition process generally involves close collaboration between a user and a system in an interactive way, in which a user would iteratively direct the system to perform specific component functions and the results from multiple steps can be synthesized to generate solutions to a complex problem. Most component-level tasks can be abstracted, studied and addressed in isolation, and evaluated using the reusable test collection methodology with reproducible experiments. While this is clearly necessary to allow for systematic progress to be made, evaluation of individual components alone is insufficient. It is arguably more important to view them as components of a larger information access system and study how to evaluate the whole system from a user’s perspective. Indeed, the ultimate goal of evaluation is to measure how well the user is aided in achieving their end goal. It is vitally important to do this evaluation correctly; if not done appropriately, it would mislead us to draw wrong conclusions or deploy an inferior application system that negatively impacts the user experience.
@@ 1.4 User Simulation
What do we exactly mean by user simulation? Informally, user simulation is to have an intelligent agent simulate how a user interacts with a system. The agent can be built based on models/algorithms/rules and any knowledge we have about the user (their behaviour, knowledge, etc.). The agent can also have parameters that can be varied in a meaningful way to simulate variations of users. Once a user simulator is constructed, it can then be used to interact with any system that needs to be evaluated. In turn, we can measure the system’s utility based on the observed behaviour of the agent while interacting with the system. Simulation thus has the potential to enable repeatable and reproducible evaluations at a low cost, without using invaluable user time (human assessor time or online experimentation bandwidth). Further, simulation can augment traditional evaluation methodologies by offering possibilities to gain insights into how system performance changes under different conditions and user behaviour.
1.4.1 Problem Definition
User simulation is the process of modeling a user’s behaviour and decision-making patterns within an interactive system, specifically designed to mimic and predict how a user will act in various interaction contexts or scenarios related to completing a task. To effectively simulate a user’s behaviour within an interactive system, configuration variables that influence this behaviour must be defined:
  • Task ($T$): A user’s behaviour varies according to nature of the user’s task. Tasks vary in complexity, and different tasks require different types and levels of interaction, decision-making processes, and completion strategies.
  • System ($S$): A user’s behaviour depends on the system they interact with. This includes the system’s functionality, user interface, and overall usability and support for task goals. It is the system that dictates the types of possible actions, denoted as $\mathcal{A}$, that a user can perform at any given point during their interactions.
  • User information ($U$): Different users may behave differently when completing the same task using the same system. Simulations must account for variations in individual user characteristics such as age, technical proficiency, preferences, and cognitive styles.
With these variables defined, the task of user simulation can be stated as the following computational problem:
Given the variables $T$, $S$, and $U$, the goal is to create an agent that can simulate every action that user $U$ may take when attempting to complete task $T$ using system $S$.
This problem involves developing a computational model that can dynamically generate user actions, reflecting the behavioural patterns and decision-making processes of a user, based on a specific task context. Formally, we define the computational model as $\pi:\mathcal{S}\rightarrow A$, where $\mathcal{S}=(T,U,S,H)$ represents the current state, encompassing information about the task $T$, system $S$, user $U$, as well as the history of previous interactions $H$ (including the actions taken by the user, the responses provided by the system, and any other relevant events that have occurred during the interaction), and $A\in\mathcal{A}$ is the action taken by the (simulated) user. The choice of computational model (e.g., rule-based, probabilistic, or machine-learned algorithm) is influenced by the nature of the task, system, and user information.
1.4.2 Scope
User simulation encompasses a wide spectrum, ranging from predicting single actions to modeling complex behaviour across multiple tasks. In our formulation, this scope is primarily determined by how the task information ($T$) is defined. At one end of the spectrum, $T$ might represent a very specific interaction context, such as predicting whether a user would click on a particular search result snippet. Here, the focus is on simulating a single, isolated action. Moving along the spectrum, $T$ could encompass a sequence of actions within a given context, such as reformulating search queries within a search session, requiring the model to consider dependencies between actions. Further expanding the scope, $T$ might represent an entire task, such as finding information on a particular topic or completing a purchase, where the simulation would involve multiple sequences of interactions. Finally, at the broadest level, $T$ could encompass a user’s general preferences and behaviour across various tasks, necessitating models that capture long-term patterns and adapt to different contexts. Thus, by varying the granularity and breadth of $T$, our formulation allows for user simulations in a wide range of application scenarios at different levels of complexity. Table 1.1 lists specific examples of user simulation for various information access tasks.
Table 1.1: Examples of user simulation, ranging from single actions to more complex behaviours.
Task ($T$) System ($S$) User information ($S$) Actions ($\mathcal{A}$)
Rating a product to express satisfaction E-commerce website with product pages and rating features User’s purchase history, browsing behaviour, and demographic information Browsing, Rating
Refining a search query to find specific information Search engine with a search box, query suggestions, and navigable search result lists User’s initial query, search history, and click behaviour Reformulating, Clicking
Collecting as many relevant information items as possible Search engine with a query box and navigable search result lists University researcher conducting a comprehensive literature review on a topic Querying, Clicking
Finding a movie to watch Recommender system with slates of items Previous watch history Clicking, Watching
Seeking assistance with a technical issue Conversational assistant with natural language chat interface User’s description of the problem, technical expertise, and previous interactions Prompting
1.4.3 Approaches
Approaches to user simulation can be classified based on the specific types of actions that they attempt to simulate. For example, some approaches may simulate how a user generates a query while others might simulate how a user responds to a search result list (e.g., simulating when a user might click on or skip a result). Approaches simulating different types of actions can also be combined to simulate a whole session of actions of a user. As will be elaborated later in the book, most existing work tends to focus on simulating each type of action separately with significantly less work on simulating a whole user session.
The problem of simulating a user’s action can often be framed as a classification problem when there is a relatively small set of actions to choose from; for example, simulation of a user’s clicking action may be framed as a binary classification problem, where the algorithm would predict whether the simulated user would click on a result or not after examining a snippet. When there are potentially infinitely many actions to choose from (e.g., when formulating a query, any valid query would be potentially an option), in practice, we often make assumptions to restrict the number of actions to be considered when simulating those actions (e.g., limit the length of a query to be considered).
With the problem framed as a classification problem, different approaches generally vary in how they perform the classification (equivalently prediction) task. At a high level, we can distinguish two broad approaches: model-based and data-driven.
  • Model-based approaches may be based on rules designed with knowledge about how users behave or on interpretable probablistic models that can more flexibly capture uncertainties using interpretable parameters. The parameters of such models may be set heuristically or empirically derived from observed user data. By varying those parameters, different types of users can be simulated.
  • Data-driven (or machine-learned) approaches emphasize maximizing accuracy of fitting any observed real user data, without necessarily imposing interpretability. Almost all such approaches are based on supervised machine learning, notably using deep neural networks which can learn effective, but non-interpretable representations from the data for predictive modeling.
These two families of approaches may also be combined, e.g., by utilizing model-based techniques to compute effective features for data-driven approaches or employing machine-learned models in specific components of model-based approaches. However, interpretability is desirable when building user simulators for evaluation to ensure that evaluation results are meaningful and to allow for the testing of verifiable hypotheses. Hence, this book primarily focuses on interpretable model-based approaches.
1.4.4 Uses of Simulation
In general, user simulation has many uses, including at least the following:
@@ 1.4.4 Uses of Simulation
  • Augmenting data with human feedback and generating synthetic data with the purpose of training machine learning models and addressing data scarcity or privacy concerns. More broadly, user simulation can facilitate machine learning approaches that require human input (interactive learning, reinforcement learning, or human-in-the-loop systems).
We note that all these uses require similar techniques, but our focus is on evaluation.
1.4.5 Requirements and Desiderata
When utilizing user simulation for system evaluation, it is critical that simulators provide reliable and insightful assessments. Two essential properties that ensure this are validity and interpretability.
  • Validity: Simulated users must exhibit behaviours that align with empirical observations of real user behaviour in similar contexts. This includes both high-level strategies (e.g., information seeking patterns) and low-level actions (e.g., clicking behaviour). Without validity, the insights gained from simulation cannot be trusted.
  • Interpretability: While not strictly a requirement, interpretability is a highly desirable property. Interpretability means that the simulated behaviour can be understood and adjusted through controllable parameters. This allows researchers to (1) understand why the simulator produced certain behaviours and (2) investigate how changes in specific parameters influence the behaviour of users. In general, as user behaviour and preferences vary significantly across users, interpretability is needed to facilitate interpretation of the evaluation results generated by user simulation, i.e., to understand what kind of real world users can be expected to produce similar results.
However, while striving for high validity is important, simulation does not need to be perfect in order to be useful. For example, although the relevance judgments in almost all the test collections for information retrieval evaluation are incomplete, the conclusions about relative performance of different retrieval systems tend not be affected much by the approximation made in a test collection [Voorhees:2000:IPM]. In fact, creating a “perfect” user simulator, i.e., one that flawlessly replicates human behaviour across all possible tasks and contexts, is likely an AI-complete problem, on par with achieving Artificial General Intelligence (AGI, cf. Section 9.4). When evaluating systems, we are often interested in a relative comparison between them with regards to some measure of utility, which is a weaker requirement than quantifying the actual utility of technology (in terms of some measurable impact, such as enhanced productivity or user satisfaction). Nevertheless, the practical utility of user simulation for relative comparisons lies in its sensitivity: the better an evaluation can distinguish between systems, the more practically useful it is.
While both validity and interpretability are desirable, there often exists a trade-off between the two. Data-driven (machine-learned) simulators, trained on large datasets of real user behaviour, can often achieve high predictive accuracy, capturing complex patterns and nuances in user actions. However, this predictive power comes at the cost of reduced interpretability. The internal workings of these models, often involving complex neural networks or ensemble methods, can be opaque and difficult to understand. This makes it challenging to pinpoint the specific reasons behind a simulated user’s behaviour or to adjust the model’s parameters in a controlled manner.
Beyond the essential requirements of validity and interpretability, there are several other desirable properties that can enhance the realism of user simulation.
  • Cognitive plausibility: The decision-making processes underlying simulated user behaviour should be grounded in theories or models of human cognition, ensuring that the simulated actions are not arbitrary or random.
  • Variation: While reflecting general user behaviour patterns, simulated users should also exhibit variability and occasional outliers, “not replicating average behaviour completely” [Bignold:2021:Biomimetics]. That is, simulation should reflect the unpredictable nature of real human interactions.
  • Adaptability: Simulated users should be able to learn from their interactions with the system, update their expectations about the system and adjust their behaviour accordingly [Balog:2021:DESIRES].
By incorporating these desirable properties, user simulators can achieve a higher level of realism and sophistication, enabling more accurate predictions of user behaviour and more insightful evaluations of interactive systems.